how do i do -2x+4y=0
-8x+y=15

Answer by nabla(475)

Multiply the first equation by -4. You get
8x-16y=0
You leave the second equation intact:
-8x+y=15
Now, add the equations:
-15y=15
This implies y=-1.
Put y=-1 into the first equation to get -2x-4=0 this implies x=-2.
You can verify that (-2,-1) works by substituting into the second equation:
16-1=15; 15=15.